OpenStudy (anonymous):

if lim(x-->0)f(x)=-1find the lim(x-->0) (f(x)+1)sin(pi/x^2)

OpenStudy (jtvatsim):

Could we use the squeeze theorem on this one?

OpenStudy (jtvatsim):

Consider that \[(f(x) + 1) \cdot -1 \leq (f(x) + 1) \cdot \sin(\pi/x^2) \leq (f(x) + 1) \cdot 1\] hence, \[\lim_{x\rightarrow0}(f(x) + 1) \cdot -1 \leq \lim_{x\rightarrow0}(f(x) + 1) \cdot \sin(\pi/x^2) \leq \lim_{x\rightarrow0}(f(x) + 1) \cdot 1\] that is, \[ 0\leq \lim_{x\rightarrow0}(f(x)+1) \cdot \sin(\pi/x^2) \leq 0\] and our desired limit is simply 0.

OpenStudy (jtvatsim):

The sin(pi/x^2) feels uncomfortable since we "know" that the inside is heading toward something undefined (or infinite), but the sin function itself is always stuck between -1 and 1 no matter what the inside is.